# FedNoisy: Federated Noisy Label Learning Benchmark

### v1

@@ Abstract
Federated learning has gained popularity for distributed learning without aggregating sensitive data from clients. But meanwhile, the distributed and isolated nature of data isolation may be complicated by data quality, making it more vulnerable to noisy labels. Many efforts exist to defend against the negative impacts of noisy labels in centralized or federated settings. However, there is a lack of a benchmark that comprehensively considers the impact of noisy labels in a wide variety of typical FL settings. In this work, we serve the first standardized benchmark that can help researchers fully explore potential federated noisy settings. Also, we conduct comprehensive experiments to explore the characteristics of these data settings and unravel challenging scenarios on the federated noisy label learning, which may guide method development in the future. We highlight the 20 basic settings for more than 5 datasets proposed in our benchmark and standardized simulation pipeline for federated noisy label learning. We hope this benchmark can facilitate idea verification in federated learning with noisy labels. FedNoisy is available at https://github.com/SMILELab-FL/FedNoisy.
1 Introduction
@@
Federated learning (FL) is a flourishing machine learning research area in recent years, which enables collaborative learning among different data centers (e.g., medical institutes and financial organizations) or small data collectors (e.g., mobile phones and smart wearable devices) to achieve promising model performance under restrictions on data sharing. Given the benefits of privacy protection and the efficiency of knowledge sharing, FL has attracted considerable attention and has been applied in various application fields, such as image classification [1, 2, 3, 4], FinTech [5, 6], text classification [7, 8], graph classification [9], human activity recognition [10, 11], and healthcare [12, 13, 14].
@@
A common yet implicit assumption in most existing FL research is that the local dataset on each client is carefully annotated with high-quality labels. However, in many real-world settings, well-labeled datasets are expensive, time-consuming, and can be prohibitive to acquire [15], for example, in medical applications [16]. Also, there may exist malicious attackers in FL systems that use label-flipping attacks [17] to intentionally manipulate local samples. In applications with low data quality, omitting clients with unreliable samples may significantly reduce training data, while manually cleaning data can induce prohibitive costs.
@@
It is well-studied that noisy labels often lead to poor model accuracy and degraded generalization capability [18]. The non-IID distribution in FL settings, as a result of intra-client class imbalance and inter-client heterogeneous distribution, may further complicate the noisy label and aggravate the negative impacts of noisy labels in learning. This complication makes existing centralized noisy label learning (CNLL) less effective and requires joint consideration of noisy label learning (NLL) and heterogeneity in FL.
@@
Therefore, federated noisy label learning (FNLL) emerges to fill the urgent need to utilize data with noisy labels in the FL training process. There are many recent efforts towards FNLL. However, they are evaluated under different experimental settings and are thus hard to comprehensively compare and choose in practical applications, due to the lack of standard benchmarks and pipelines. The differences in experimental settings are due to the following factors. 1) Datasets. Existing works on FNLL are evaluated on different datasets. For synthetic noise, [19, 20, 21] evaluate on CIFAR-10 which is common in both FL and CNLL; [22] uses PD-Tremor dataset and [23] uses Adult [24], and HFP [25] datasets, which are rarely used in CNLL [15]. 2) Data partition schemes. Though most works [26, 27, 28] use common IID partition schemes, they have differences in non-IID treatments and therefore different data partitions. [21] used Dirichlet partition, which is commonly used among FL works. [22] used real-world partition generated by hospitals. [29] deployed an extreme distribution scheme that allows only a single class on each client. 3) Types of label noise. Existing FNLL works were studied under a variety of generation processes of label noise in the FL environment. [19, 21] evaluated both symmetric noise and asymmetric label noise as in centralized NLL research. [20] only evaluated symmetric noise; [29] evaluated open-set noise and [30] used other types of noise. When it comes to noise distribution on different clients, different works introduced different settings: [22, 29, 27] allowed clients to be either clean or noised, with all noised clients contain the same level of label noise; [21] separated clients into 4 groups with 4 different noise levels; [23] set all clients with same noise ratio. 4) Limited baselines. Most existing FNLL works combine only SOTA CNLL methods with FL algorithms as baselines in experiments, partially due to the fact that the implementations of FNLL approaches did not conform to a standard benchmark/pipeline, which makes consistent comparisons across FNLL hardly possible. For example, [20] combined DivideMix[31]/JointOpt[32] with FedAvg as baselines; [19] deployed Co-teaching on FedAvg as baseline. Most of these methods belong to limited categories in CNLL, i.e., sample selection methods [15], while dismissing others in CNLL. These limited baselines could lead to barriers when diving deep into FNLL, without fully exploring the potential capability of existing NLL methods. We attribute the scarcity of baselines to the lack of a well-designed FL framework, which includes a variety of datasets, easy access to different partition schemes, ready-to-go simulation environments, and extensible baselines covering a wide scope of methodologies.
@@
  • We implement an FNLL codebase with ready-to-go federated noisy-label settings. The codebase comprises more than 8 FNLL baselines covering a broad scope of existing CNLL research categories. All baselines are presented in a unified FL workflow, capable of further improvement with different FL algorithms in a pluggable manner. We have included 5 datasets with built-in interfaces for the convenience of evaluating existing methods.
  • We conduct a comprehensive evaluation following a standardized experimental pipeline to show the intriguing characteristics of FNLL data settings. We have several major findings: Unlike the case in a centralized setting, symmetric noise can be harder than asymmetric ones under some Non-IID partitions; Also, focusing on denoising is profitable when the noise ratio is small, while handling Non-IID data is more lucrative given large noise ratio.
2 Related Works
@@
Noisy Label Learning. Noisy label learning (NLL) aims to train robust DNNs with noisy labels. Different approaches have been proposed to address this problem in the centralized learning setting, and they can be roughly divided into five categories [15]: i) robust architecture [33, 34]; ii) robust regularization methods [35, 36], iii) robust loss function design [37, 38], iv) loss adjustment methods [39, 40], and v) sample selection methods [31, 41]. We refer readers to related NLL survey papers for more details (e.g., [15]). In FedNoisy, we provide at least one method from each of the last four categories, combining them with FL algorithms as baselines for FNLL research. We plan to support more methods from different categories of NLL in the future.
@@
Federated Noisy Label Learning. Despite diverse research devoted to Non-IID problems and communication efficiencies, most assume that the training datasets only contain clean labels. Some works have been trying to alleviate the label noise in FL recently, which mainly follow three lines: i) client selection, ii) client reweighting, and iii) sample selection. Client selection methods [30, 21, 20] select clients with high-quality labels based on different evaluation metrics and only allow selected clients to participate FL training rounds. Client reweighting methods [42, 22, 20] evaluate each client based on the quality of local datasets and assign lower weighting scores to noisy clients during aggregation without abandoning them. Sample selection methods evaluate each sample on clients, for example, per-sample loss, and exclude noisy samples for later training [20, 29, 19]. In addition to the previous mainstream approaches, researchers will reuse the low-quality samples via pseudo-labeling for further performance improvement [20, 19, 21].
@@
Federated Learning Benchmark. Standardized benchmark empowers researchers to follow previous research with less effort and facilitates the validation of research ideas. There are several benchmark works in the field of FL, including standardized simulation environments [43, 44, 45], heterogeneous data distribution [46, 47], and specific applications [13, 48, 49]. However, to the best of our knowledge, no existing benchmark is proposed for FL with noisy labels, which leads to the non-standardized simulation pipeline and unfair comparison among related research. The proposed FedNoisy bridges the gap between the FL community and existing CNLL research and provides a roadmap for the experiment setting for future FNLL works.
3 Problem Formulation
@@
Federated Learning (FL) collaboratively learns a model among multiple participants without directed data sharing [50]. Here we consider a representation FL setting: $C$-class classification in a horizontal federated learning system with a central server. Let $\mathcal{D}=\{(x^{n},y^{n})\}_{n=1}^{N}$ denote the global training dataset, where $x^{n}\in\mathcal{X}$ denotes a sample, and $y^{n}\in\mathcal{Y}=[C]$ is its ground-truth label. Suppose there are $K$ clients in the federated system splitting $\mathcal{D}$ without overlap, and for each client $k$ ($k=1,\ldots,K$), there is a local dataset $\mathcal{D}_{k}=\{(x_{k}^{n},y_{k}^{n})\}_{n=1}^{N_{k}}$ such that $N=\sum_{k=1}^{K}{N_{k}}$. In typical FL problems, the distribution of $\mathcal{D}_{k}$ varies for different clients and is called Non-IID.
@@
$\min_{w}\mathcal{G}(F_{1}(w),\ldots,F_{K}(w)),$ (1)
where $F_{k}(w)$ is the local objective for client $k$, and $\mathcal{G}$ is a function that combines local objectives together. $\mathcal{G}(\cdot)$ can be simple weighted average of $\{F_{k}(w)\}_{k\in[K]}$ as in FedAvg [50]. Local objective $F_{k}(\cdot)$ can be the empirical risk on the local dataset $\mathcal{D}_{k}$:
$F_{k}(w)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\left[\ell(y,f_{k}(x;w))\right],$ (2)
@@
Noisy Label Learning aims to mitigate the performance drops caused by noised training labels. The poor label qualities often arise from limited annotation resources where many labels are not fully examined in practice. Given a well-labeled dataset, the corrupted training dataset is forged by randomly changing true labels following a predefined distribution and then a model is trained on it and evaluated on a clean test set. The corrupted label $\tilde{y}$ hinges on the ground-truth label $y$ while independent of its sample $x$. The noise transition matrix $\mathbf{T}\in[0,1]^{C\times C}$ denotes all possible flipping probability between classes, that is, $\mathbf{T}_{ij}=P(\tilde{y}=j\mid y=i)$. Given a constant noise ratio $\varepsilon\in[0,1]$, representative instantiations of such noise could symmetric noise or asymmetric one. (1) In the symmetric setting, clean labels are uniformly corrupted into other labels with equal probability, i.e., $\forall_{i\neq j}\mathbf{T}_{ij}=\frac{\varepsilon}{C-1}$ and $\forall_{i=j}\mathbf{T}_{ij}=1-\varepsilon$. (2) In the asymmetric setting, the label will be only flipped to a particular label, i.e., $\forall_{i=j}\mathbf{T}_{ij}=1-\varepsilon$, $\exists_{c\neq i}\mathbf{T}_{ic}=\varepsilon$, and $\forall_{j\neq i,j\neq c}\mathbf{T}_{ij}=0$. Such asymmetric noise is also called pairwise flipping. In some literature [51], the corruption probability can be formulated as $P(\tilde{y}=j\mid y=i,x)$, namely instance-dependent noise. In this paper, we only consider symmetric noise, asymmetric noise following common practice, and also include dataset with real-world noise.
@@
$F_{k}(w)=\mathbb{E}_{(x,\tilde{y})\sim\tilde{\mathcal{D}}_{k}}\left[\ell(\tilde{y},f_{k}(x;w))\right],$ (3)
@@ 4 Benchmark Design
4.1 Label-Noise Simulations for Heterogeneous Clients
@@
Figure 1: Federated noise scenes provided in FedNoisy. Left: globalized noise; middle: localized noise; right: real-word noise. Globalized noise follows a constant noise ratio $\varepsilon_{global}$ in label corruption, while localized noise draws localized noise ratio $\varepsilon_{k}$ from a uniform distribution on each client $k$.
@@
Localized noise scene. In real-world federated systems, label noise may be attributed to clients’ local data collection process, which means each client has an independent label corruption process with local label noise distribution. For example, hospitals with limited medical resources are more likely to have higher misdiagnosis rates [52]. To simulate this, we propose localized noise scene, in which different clients have different local noise ratios of noisy label distribution. And it can also be called heterogeneous label noise.
@@
In localized noise scene, a clean global dataset $\mathcal{D}$ is first split into $K$ local clean datasets $\{\mathcal{D}_{1},\ldots,\mathcal{D}_{K}\}$, following IID/non-IID fashion. For each client $k$, the local noise distribution is synthesized using a local noise transformation matrix $\mathbf{T}_{k}\in[0,1]^{C_{k}\times C_{k}}$ generated from local noise ratio $\varepsilon_{k}$, where $\varepsilon_{k}$ is drawn from a uniform distribution $\mathcal{U}(\varepsilon_{min},\varepsilon_{max})$ with $\varepsilon_{min},\varepsilon_{max}\in[0,1]$, and $C_{k}$ is the total number of classes on clean local dataset $\mathcal{D}_{k}$. We flip sample labels of the local clean dataset $\mathcal{D}_{k}$ using $\mathbf{T}_{k}$ for symmetric/asymmetric noise and obtain noisy local dataset $\tilde{\mathcal{D}}_{k}$ for each client $k$. Also, the label noising process on each client only allows label flipping among classes in the current local client. To be more concrete, given an FL system with samples labels of $\{1,2,3,4,5\}$ in the global dataset, if a particular client only holds a local dataset with class label $\{1,3,5\}$, then there are no chances for the existence of noisy samples with class label $\{2,4\}$ on this client after label corruption. This constraint of the localized scene closely resembles many real-world cases: a dentist may misdiagnose different dental diseases while unlikely to diagnose lung cancer. The middle part of Figure 1 shows the generation process of the localized noise scene.
@@
Real-world noise scene. Unlike globalized and localized noise scenes that only contain synthetic label noise, real-world noise scene evaluates against datasets inherently with mislabeled samples. Here we use one popular real-world noisy dataset in CNLL studies [15], called Clothing1M [51], as summarized in Table 1. In the real-world noise scene, we consider directly partitioning the noisy global dataset $\tilde{\mathcal{D}}$ into $\{\tilde{\mathcal{D}}_{1},\ldots,\tilde{\mathcal{D}}_{K}\}$ via IID/non-IID fashion to create an FL setting with real-world label noise. The right part of Figure 1 shows the generation process of the real-world noise scene.
4.2 Datasets and Data Heterogeneity Simulation

### v4

@@ Abstract.
Federated learning has gained popularity for distributed learning without aggregating sensitive data from clients. But meanwhile, the distributed and isolated nature of data isolation may be complicated by data quality, making it more vulnerable to noisy labels. Many efforts exist to defend against the negative impacts of noisy labels in centralized or federated settings. However, there is a lack of a benchmark that comprehensively considers the impact of noisy labels in a wide variety of typical FL settings. In this work, we serve the first standardized benchmark that can help researchers fully explore potential federated noisy settings. Also, we conduct comprehensive experiments to explore the characteristics of these data settings and the comparison across baselines, which may guide method development in the future. We highlight the 20 basic settings for 6 datasets proposed in our benchmark and standardized simulation pipeline for federated noisy label learning, including implementations of 9 baselines. We hope this benchmark can facilitate idea verification in federated learning with noisy labels. FedNoisy is available at https://github.com/SMILELab-FL/FedNoisy.
Noisy Label, Federated Learning, Benchmark
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Distributed algorithmsccs: Computing methodologies Machine learning algorithmsccs: Information systems Data cleaning
1. Introduction
@@
Federated learning (FL) is a flourishing machine learning research area in recent years, which enables collaborative learning among different data centers (e.g., medical institutes and financial organizations) or small data collectors (e.g., mobile phones and smart wearable devices) to achieve promising model performance under restrictions on data sharing. Given the benefits of privacy protection and the efficiency of knowledge sharing, FL has attracted considerable attention and has been applied in various application fields, such as image classification (Jeong et al., 2021; Li et al., [n. d.]; Zhang et al., 2021; Zhou et al., 2022), FinTech (Zhang et al., 2023; Long et al., 2020), text classification (Zhang et al., 2022a, b), graph classification (He et al., 2021b), human activity recognition (Sozinov et al., 2018; Ouyang et al., 2021), and healthcare (Liu et al., 2021a; du Terrail et al., 2022; Xu et al., 2021).
@@
A common yet implicit assumption in most existing FL research is that the local dataset on each client is carefully annotated with high-quality labels. However, in many real-world settings, well-labeled datasets are expensive, time-consuming, and can be prohibitive to acquire (Song et al., 2022b), for example, in medical applications (Karimi et al., 2020). Also, there may exist malicious attackers in FL systems that use label-flipping attacks (Lyu et al., 2020) to intentionally manipulate local samples. In applications with low data quality, omitting clients with unreliable samples may significantly reduce training data, while manually cleaning data can induce prohibitive costs.
@@
It is well-studied that noisy labels often lead to poor model accuracy and degraded generalization capability (Zhang et al., 2017). The non-IID distribution in FL settings, as a result of intra-client class imbalance and inter-client heterogeneous distribution, may further complicate the noisy label and aggravate the negative impacts of noisy labels in learning. This complication makes existing centralized noisy label learning (CNLL) less effective and requires joint consideration of noisy label learning (NLL) and heterogeneity in FL.
@@
Therefore, federated noisy label learning (FNLL) emerges to fill the urgent need to utilize data with noisy labels in the FL training process. There are many recent efforts towards FNLL. However, they are evaluated under different experimental settings and are thus hard to comprehensively compare and choose in practical applications, due to the lack of standard benchmarks and pipelines. The differences in experimental settings are due to the following factors. 1) Datasets. Existing works on FNLL are evaluated on different datasets. For synthetic noise, (Yang et al., 2022; Xu et al., 2022; Wang et al., 2022) evaluate on CIFAR-10 which is common in both FL and CNLL; (Chen et al., 2020) uses PD-Tremor dataset and (Duan et al., 2022) uses Adult (Kohavi et al., 1996), and HFP (Chicco and Jurman, 2020) datasets, which are rarely used in CNLL (Song et al., 2022b). 2) Data partition schemes. Though most works (Li et al., 2021; Tahmasebian et al., 2022; Li et al., 2022b) use common IID partition schemes, they have differences in non-IID treatments and therefore different data partitions. (Wang et al., 2022) used Dirichlet partition, which is commonly used among FL works. (Chen et al., 2020) used real-world partition generated by hospitals. (Tuor et al., 2020) deployed an extreme distribution scheme that allows only a single class on each client. 3) Types of label noise. Existing FNLL works were studied under a variety of generation processes of label noise in the FL environment. (Yang et al., 2022; Wang et al., 2022) evaluated both symmetric noise and asymmetric label noise as in centralized NLL research. (Xu et al., 2022) only evaluated symmetric noise; (Tuor et al., 2020) evaluated open-set noise and (Wang et al., 2020a) used other types of noise. When it comes to noise distribution on different clients, different works introduced different settings: (Chen et al., 2020; Tuor et al., 2020; Tahmasebian et al., 2022) allowed clients to be either clean or noised, with all noised clients contain the same level of label noise; (Wang et al., 2022) separated clients into 4 groups with 4 different noise levels; (Duan et al., 2022) set all clients with same noise ratio. 4) Limited baselines. Most existing FNLL works combine only SOTA CNLL methods with FL algorithms as baselines in experiments, partially due to the fact that the implementations of FNLL approaches did not conform to a standard benchmark/pipeline, which makes consistent comparisons across FNLL hardly possible. For example, (Xu et al., 2022) combined DivideMix(Li et al., 2020)/JointOpt(Tanaka et al., 2018) with FedAvg as baselines; (Yang et al., 2022) deployed Co-teaching on FedAvg as baseline. Most of these methods belong to limited categories in CNLL, i.e., sample selection methods (Song et al., 2022b), while dismissing others in CNLL. These limited baselines could lead to barriers when diving deep into FNLL, without fully exploring the potential capability of existing NLL methods. We attribute the scarcity of baselines to the lack of a well-designed FL framework, which includes a variety of datasets, easy access to different partition schemes, ready-to-go simulation environments, and extensible baselines covering a wide scope of methodologies.
@@
  • We implement an FNLL codebase with ready-to-go federated noisy-label settings. The codebase comprises 9 FNLL baselines covering a broad scope of existing CNLL research categories as well as the most recent FNLL work. All baselines are presented in a unified FL workflow, capable of further improvement with different FL algorithms in a pluggable manner. We have included 6 datasets with built-in interfaces for the convenience of evaluating existing methods.
  • We conduct a comprehensive evaluation following a standardized experimental pipeline to show the intriguing characteristics of FNLL data settings as well as the comparison between the included FNLL baselines. Our findings highlight several key aspects: Evaluating the generalizability of FNLL algorithms across diverse model architectures is essential; FNLL algorithms need to be robust across different datasets and noise settings; SCE and GCE emerge as the best baselines in balancing accuracy and computational efficiency, recommending their use as primary baselines in FNLL algorithm development; Unlike the case in a centralized setting, symmetric noise can be harder than asymmetric ones under some Non-IID partitions; While focusing on denoising proves beneficial at low noise ratios, addressing Non-IID data becomes more critical when dealing with higher noise levels.
2. Related Works
@@
Noisy Label Learning. Noisy label learning (NLL) aims to train robust DNNs with noisy labels. Different approaches have been proposed to address this problem in the centralized learning setting, and they can be roughly divided into five categories (Song et al., 2022b): i) robust architecture (Yao et al., 2019; Lee et al., 2019); ii) robust regularization methods (Xia et al., 2021; Zhang et al., 2018), iii) robust loss function design (Wang et al., 2019; Ma et al., 2020), iv) loss adjustment methods (Patrini et al., 2017; Arazo et al., 2019), and v) sample selection methods (Li et al., 2020; Han et al., 2018). We refer readers to related NLL survey papers for more details (e.g., (Song et al., 2022b)). In FedNoisy, we provide at least one method from each of the last four categories, combining them with FL algorithms as baselines for FNLL research. We plan to support more methods from different categories of NLL in the future.
@@
Federated Noisy Label Learning. Despite diverse research devoted to Non-IID problems and communication efficiencies, most assume that the training datasets only contain clean labels. Some works have been trying to alleviate the label noise in FL recently, which mainly follow three lines: i) client selection, ii) client reweighting, and iii) sample selection. Client selection methods (Wang et al., 2020a, 2022; Xu et al., 2022) select clients with high-quality labels based on different evaluation metrics and only allow selected clients to participate FL training rounds. Client reweighting methods (Fang and Ye, 2022; Chen et al., 2020; Xu et al., 2022) evaluate each client based on the quality of local datasets and assign lower weighting scores to noisy clients during aggregation without abandoning them. Sample selection methods evaluate each sample on clients, for example, per-sample loss, and exclude noisy samples for later training (Xu et al., 2022; Tuor et al., 2020; Yang et al., 2022). In addition to the previous mainstream approaches, researchers will reuse the low-quality samples via pseudo-labeling for further performance improvement (Xu et al., 2022; Yang et al., 2022; Wang et al., 2022).
@@
Federated Learning Benchmark. Standardized benchmark empowers researchers to follow previous research with less effort and facilitates the validation of research ideas. There are several benchmark works in the field of FL, including standardized simulation environments (Beutel et al., 2020; He et al., 2020; Zeng et al., 2023), heterogeneous data distribution (Li et al., 2022a; Caldas et al., 2018), and specific applications (du Terrail et al., 2022; Song et al., 2022a; He et al., 2021a). However, to the best of our knowledge, no existing benchmark is proposed for FL with noisy labels, which leads to the non-standardized simulation pipeline and unfair comparison among related research. The proposed FedNoisy bridges the gap between the FL community and existing CNLL research and provides a roadmap for the experiment setting for future FNLL works.
3. Problem Formulation
@@
Federated Learning (FL) collaboratively learns a model among multiple participants without directed data sharing (McMahan et al., 2017). Here we consider a representation FL setting: $C$-class classification in a horizontal federated learning system with a central server. Let $\mathcal{D}=\{(x^{n},y^{n})\}_{n=1}^{N}$ denote the global training dataset, where $x^{n}\in\mathcal{X}$ denotes a sample, and $y^{n}\in\mathcal{Y}=[C]$ is its ground-truth label. Suppose there are $K$ clients in the federated system splitting $\mathcal{D}$ without overlap, and for each client $k$ ($k=1,\ldots,K$), there is a local dataset $\mathcal{D}_{k}=\{(x_{k}^{n},y_{k}^{n})\}_{n=1}^{N_{k}}$ such that $N=\sum_{k=1}^{K}{N_{k}}$. In typical FL problems, the distribution of $\mathcal{D}_{k}$ varies for different clients and is called Non-IID.
@@
(1) $\min_{w}\mathcal{G}(F_{1}(w),\ldots,F_{K}(w)),$
where $F_{k}(w)$ is the local objective for client $k$, and $\mathcal{G}$ is a function that combines local objectives together. $\mathcal{G}(\cdot)$ can be simple weighted average of $\{F_{k}(w)\}_{k\in[K]}$ as in FedAvg (McMahan et al., 2017). Local objective $F_{k}(\cdot)$ can be the empirical risk on the local dataset $\mathcal{D}_{k}$:
(2) $F_{k}(w)=\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\left[\ell(y,f_{k}(x;w))\right],$
@@
Noisy Label Learning aims to mitigate the performance drops caused by noised training labels. The poor label qualities often arise from limited annotation resources where many labels are not fully examined in practice. Given a well-labeled dataset, the corrupted training dataset is forged by randomly changing true labels following a predefined distribution and then a model is trained on it and evaluated on a clean test set. The corrupted label $\tilde{y}$ hinges on the ground-truth label $y$ while independent of its sample $x$. The noise transition matrix $\mathbf{T}\in[0,1]^{C\times C}$ denotes all possible flipping probability between classes, that is, $\mathbf{T}_{ij}=P(\tilde{y}=j\mid y=i)$. Given a constant noise ratio $\varepsilon\in[0,1]$, representative instantiations of such noise could symmetric noise or asymmetric one. (1) In the symmetric setting, clean labels are uniformly corrupted into other labels with equal probability, i.e., $\forall_{i\neq j}\mathbf{T}_{ij}=\frac{\varepsilon}{C-1}$ and $\forall_{i=j}\mathbf{T}_{ij}=1-\varepsilon$. (2) In the asymmetric setting, the label will be only flipped to a particular label, i.e., $\forall_{i=j}\mathbf{T}_{ij}=1-\varepsilon$, $\exists_{c\neq i}\mathbf{T}_{ic}=\varepsilon$, and $\forall_{j\neq i,j\neq c}\mathbf{T}_{ij}=0$. Such asymmetric noise is also called pairwise flipping. In some literature (Xiao et al., 2015), the corruption probability can be formulated as $P(\tilde{y}=j\mid y=i,x)$, namely instance-dependent noise. In this paper, we only consider symmetric noise, asymmetric noise following common practice, and also include dataset with real-world noise.
@@
(3) $F_{k}(w)=\mathbb{E}_{(x,\tilde{y})\sim\tilde{\mathcal{D}}_{k}}\left[\ell(\tilde{y},f_{k}(x;w))\right],$
@@ 4. Benchmark Design
4.1. Label-Noise Simulations for Heterogeneous Clients
@@
Figure 1. Federated noise scenes provided in FedNoisy. Left: globalized noise; middle: localized noise; right: real-word noise. Globalized noise follows a constant noise ratio $\varepsilon_{global}$ in label corruption, while localized noise draws localized noise ratio $\varepsilon_{k}$ from a uniform distribution on each client $k$.
@@
Localized noise scene. In real-world federated systems, label noise may be attributed to clients’ local data collection process, which means each client has an independent label corruption process with local label noise distribution. For example, hospitals with limited medical resources are more likely to have higher misdiagnosis rates (Zhang et al., 2020). To simulate this, we propose localized noise scene, in which different clients have different local noise ratios of noisy label distribution. And it can also be called heterogeneous label noise.
@@
In localized noise scene, a clean global dataset $\mathcal{D}$ is first split into $K$ local clean datasets $\{\mathcal{D}_{1},\ldots,\mathcal{D}_{K}\}$, following IID/non-IID fashion. For each client $k$, the local noise distribution is synthesized using a local noise transformation matrix $\mathbf{T}_{k}\in[0,1]^{C_{k}\times C_{k}}$ generated from local noise ratio $\varepsilon_{k}$, where $\varepsilon_{k}$ is drawn from a uniform distribution $\mathcal{U}(\varepsilon_{min},\varepsilon_{max})$ with $\varepsilon_{min},\varepsilon_{max}\in[0,1]$, and $C_{k}$ is the total number of classes on clean local dataset $\mathcal{D}_{k}$. We flip sample labels of the local clean dataset $\mathcal{D}_{k}$ using $\mathbf{T}_{k}$ for symmetric/asymmetric noise and obtain noisy local dataset $\tilde{\mathcal{D}}_{k}$ for each client $k$. Also, the label noising process on each client only allows label flipping among classes in the current local client. To be more concrete, given an FL system with sample labels of $\{1,2,3,4,5\}$ in the global dataset, if a particular client only holds a local dataset with class label $\{1,3,5\}$, then there are no chances for the existence of noisy samples with class label $\{2,4\}$ on this client after label corruption. This constraint of the localized scene closely resembles many real-world cases: a dentist may misdiagnose different dental diseases while unlikely to diagnose lung cancer. The middle part of Figure 1 shows the generation process of the localized noise scene.
@@
Real-world noise scene. Unlike globalized and localized noise scenes that only contain synthetic label noise, real-world noise scene evaluates against datasets inherently with mislabeled samples. Here we use two popular real-world noisy datasets in CNLL studies (Song et al., 2022b), that is, Clothing1M (Xiao et al., 2015) and WebVision (Li et al., 2017)  [^1] We use the first 50 classes of the Google subset from WebVision 1.0 resized version as noisy global dataset $\tilde{\mathcal{D}}$ in our benchmark., as summarized in Table 1. In the real-world noise scene, we consider directly partitioning the noisy global dataset $\tilde{\mathcal{D}}$ into $\{\tilde{\mathcal{D}}_{1},\ldots,\tilde{\mathcal{D}}_{K}\}$ via IID/non-IID fashion to create an FL setting with real-world label noise. The right part of Figure 1 shows the generation process of the real-world noise scene.
4.2. Datasets and Data Heterogeneity Simulation